Use the Gauss- Jordan method to solve each system of equations
x+3y-6z=7
2x-y+2z=0
x+y+2z=-1

Answer by stanbon(75887)

Use the Gauss- Jordan method to solve each system of equations
x+3y-6z=7
2x-y+2z=0
x+y+2z=-1
-----------------
Gauss-Jordan is a lot of arithmetic.
---
I used a matrix method to get:
x = 1
y = 1 2/3
z = -1/6
==============
Cheers,
Stan H.